OpenStudy (anonymous):

if the equation is an identity,write identity if has no solution rite no solution 6(4z+2)=3(8z+4)

hartnn (hartnn):

multiply them out like: a(b+c)=ab+ac so what is 6(4z+2) =?

OpenStudy (anonymous):

6(4z+2)=3(8z+4)=24z +12 =24z+12

hartnn (hartnn):

@xxfreshboy please do not provide complete solution

OpenStudy (anonymous):

uhm the Identify in this case is true which goes for all values.

OpenStudy (anonymous):

oo aiight just helping out...

OpenStudy (anonymous):

e.e idkkkk

hartnn (hartnn):

help them understand how to solve, by engaging them

hartnn (hartnn):

could you not do 6(4x+2) ?

OpenStudy (anonymous):

Distribute!

OpenStudy (anonymous):

do i have to divide

OpenStudy (anonymous):

or something